An empty pressure cooker of
volume 10 litres contain air at
atmospheric pressure 10^5 Pa and
temperature of 300 K. It contains a
whistle which has area if 0.1 cm^2
and weight of 100 gm. What should
be the Temprature of air inside so
that the whistle is just lifted up?
please need the steps of doing it

Grade:12

2 Answers

kshitij aggarwal
10 years ago

First let us find the pressure at which whistle will be lifted. 

        P = mg / A

          = 105 Pa.

so it will be lifted at 300K only.

 If say the pressure inside the cooker at 300K be some other value,  then use P/  T1 = P2  / T2

nimo jar
6 years ago
Pressure by whistle = Po + F / A = Po​ + mg/ A
= 105Pa + (0.1 * 10 N) / (0.1 * 104 m2)
= 2 * 105 Pa
Pressure inside cooker = 105Pa
for whistle to just lift, by gay lussacs law
Pwhistle / Tsurrounding = Pcooker / Tcooker
(2 * 105 Pa) / 300 K = 105 Pa / T
T = 600 K = 327oC